# Global Balance and Systemic Risk in Financial Correlation Networks

### v1

@@ H Pearson Spearman Pearson Spearman
Figure 5.1: Global Balance (square dots black line) and Average Correlation (circular dots red lines), S&P500 daily returns, $\Delta T=400$ and $\Delta t=30$, for the period (a) 2005-2016, (b) 2016-2020 and (c) zoom on the period 2008-2014.
We can extend the analysis to the correlations between the AMRI and the ratio ${\mathscr{R}}(G)$ in Eq. (17). These correlations are very close to the previous ones. Indeed, both $\kappa(G)$ and ${\mathscr{R}}(G)$ produce time series with very similar behaviors, making the two measures almost equivalent for application purposes. This fact can be explained by the presence of a large spectral gap in eigenvalues of the correlation matrices for each time window. In this case, it is possible to apply the approximate formula for the ratio in Eq. (20). This explains the almost-1 correlation between $\kappa(G)$ and ${\mathscr{R}}(G)$. We observe that the error between the exact value of the ratio in Eq. (19) and its approximate version in Eq. (20) is very small. Until October 2016, the mean relative error is $0.033\%$, with a maximum equal to $0.201\%$ in August, 2015. In the period from November 2016 to January 2020, the approximation error is on average $1.22\%$ with a max of $3.91\%$ in October 2017. The maximum is achieved in correspondence with a severe reduction of the spectral gap in 2017, see Fig. 5.3. We also observe that the network is perfectly balanced in the windows ending in February to May 2010, which the entire year 2009.
@@
Fig. 5.4 and Fig. 5.5 show the temporal evolution of the involved quantities on the whole time interval and on the partitioned time interval, respectively.
@@
Similarly, the correlations between the Global Balance and, for instance, the MRI for $H=2$ are $\rho_{\rm Pearson}=0.5307341$ and $\rho_{\rm Spearman}=0.7685456$, on the period 2005-2016 and $\rho_{\rm Pearson}=0.4994806$ and $\rho_{\rm Spearman}=0.9441659$, on the period 2016-2020. The correlation between the Global Balance and the Cumulative Risk Fraction are $\rho_{\rm Pearson}=0.9096708$ and $\rho_{\rm Spearman}=0.9203074$, on the period 2005-2016 and $\rho_{\rm Pearson}=0.9297902$ and $\rho_{\rm Spearman}=0.991976$, on the period 2016-2020. To test the robustness of our findings, we repeated the exercise over many different subsets of 50 assets randomly chosen from the original index, obtaining similar results.
Figure 5.4: Global Balance (square dots black line) and Average Correlation (circular dots red lines), a sample of $50$ assets from S&P500 daily returns, $\Delta T=100$ and $\Delta t=30$, for the period (a) 2005-2016, (b) 2016-2020 and (c) zoom on the period 2008-2014.
Figure 5.5: Time evolution of the global balance (black square dot line) and of the market rank indicator for $p=3$ and $H=2$ (blue kite dot line), $H=5$ (red triangle dot line) and $H=10$ (green circle dot line) for the period (a) 2005-2016 and (b) 2016-2020. The figures refer to a subset of $50$ assets.
@@ 5.2 Effectiveness of the Global Balance as a systemic risk measure
We perform now an empirical analysis to assess the effectiveness of the Global Balance in detecting systemic events. In order to identify systemic events, we refer to the definition in [54]. A systemic risk event occurs if the average return of all the $N$ securities falls below a threshold $\tau$ for $\Delta T$ consecutive days. We adopt a threshold equal to $\tau=-0.01$ or $\tau=-0.005$ and sliding windows of $\Delta T=20$ days.
Figure 5.6: Panel (a): Average returns of the S&P500 database in the period 2005-2020. The average returns of each asset have been computed over a $20$-days-wide sliding window with $1$-day step. The blue line represents the mean over all the assets in the same window of the average returns. The green and the red lines represent the maximum and the minimum, respectively, of the average returns in the same window. Panel (b) focuses on the mean of the average returns and highlight crisis events when the line goes below the two possible threshold represented by the two horizontal dashed lines.
In Fig, 5.6, panel (a), we plot the average returns of the S&P500 database from January 4, 2005 to March 18, 2020. The average returns of each asset have been computed over a $20$-days-wide sliding window with $1$-day step. The blue line represents the mean of the average returns over all the $385$ assets in the same window. The green and the red lines represent the maximum and the minimum value, respectively, of the average returns in the same window. Panel (b) focuses on the mean of the average returns and highlight crisis events when the line goes below the two possible thresholds represented by the two horizontal dashed lines.
In this way, we can clearly identify systemic crisis events when the mean of the average returns over all the asses goes below $\tau$, for both $\tau=-0.005$ and $\tau=-0.010$. In particular, the crisis events below $\tau=-0.010$ ranges in the time intervals September-November 2008, February 2009, July 2011, October 2019.
We study the joint distribution of the average returns and of the global balance in the sliding windows. The range of the indicator $\kappa(G)$ has been divided into $n$ bins $B_{1},B_{2},\dots,B_{n}$ and the conditional distributions of the average returns $P\left(\langle X_{it}\rangle|\,\kappa(G)\in B_{k}\right)$ have been computed on each bin. Specifically, in a first experiment the interval $[0,1]$ is divided into $5$ bins of equal width while, in a second scenario, the five intervals are $(0,0.5]$, $(0.5,0.8]$, $(0.8,0.9]$, $(0.9,0.99]$ and $(0.99,1.00]$. The necessity of testing what happens with intervals of different lengths, very small when close to 1, is a peculiarity in the framework of systemic risk. Since systemic risk are, hopefully, rare compared to the periods in which the market behaves normally, a systemic risk indicator should discriminate a few events within a multitude.
@@
Figure 5.7 displays the densities of the average returns conditioned to the values of $\kappa(G)$, obtained by using a Gaussian kernel smoothing for the whole dataset of $385$ assets (panels (a) and (b) for the weighted version of the correlation networks and panels (c) and (d) for the binary version). When $\kappa(G)$ increases, the densities move leftwards and flatten. More precisely, three effects can be recognized in relation to the increase of $\kappa(G)$: the average of the conditional distribution decreases; the standard deviation increases; the VaR [^5] The Value at Risk, $p$ VaR, measures the loss that will not be exceeded with a given confidence probability $p$ over a certain time horizon (see, for instance, [55]). We adopt $p=0.05$. of the average returns increases.
Fig. 5.8 shows the boxplots of the same distributions conditioned on the two different partitions of the $[0,1]$ interval for the global balance in the weighted and binary version. The mean of the average returns within each interval decreases as $\kappa(G)$ increases, while the standard deviation increases. This is particularly noticeable by considering the extreme intervals in Fig. 5.8, panels (b) and (d), which refer to the events in the interval $\kappa(G)\in(0.99,1.00]$.
The values of the mean of the conditional distributions in Fig. 5.8 are collected in Table 5.2, while Table 5.3 reports the values of the means computed for the smoothed distributions in Fig. 5.7. Both are characterized by a strictly negative monotonic behavior as $\kappa(G)$ increases.
@@
Finally, we computed the VaR within each band. For instance, for the whole network in the weighted version and with intervals of equal width the VaR’s are: $0.02384863$ for $0<\kappa(G)\leq 0.2$; $0.0281525$, for $0.2<\kappa(G)\leq 0.4$; $0.03010301$, for $0.4<\kappa(G)\leq 0.6$; $0.03521984$, for $0.6<\kappa(G)\leq 0.8$; and $0.05100622$ for $0.8<\kappa(G)\leq 1.0$.
This evidence shows that the global balance is a meaningful systemic risk measure. Indeed, a large value of $\kappa(G)$ signals that the market is experiencing worse performance and that the probability of big losses is high.
Finally, we check the robustness of our findings using a different data set. We refer to the constituents of the eurostoxx 50 index (ESX50). The cleaned dataset contains the daily returns of $42$ stocks from January 5, 2005 to September 17, 2020. We perform the test over windows of $45$ days. In Fig. 5.9 we plot the mean, maximum and minimum average returns of the constituents, panel (a), and a zoom on the mean, panel (b). Due to the larger size of the time windows, we detect less but more severe events. In this new European dataset, we specifically identify the systemic crisis associated with COVID-19 in January 2020 with a threshold $\tau=-0.010$. Similarly to the previous database, Fig. 5.10 and Table 5.4 illustrate the conditional smoothed distributions, the boxplots and the corresponding mean values for the weighted version of the correlation networks. We do not show the results for the binary network for the present example. The numerical results confirm the previous findings, reinforcing the interpretation of the global balance as an effective systemic risk indicator.

### v2

@@ H Pearson Spearman Pearson Spearman
We can extend the analysis to the correlations between the AMRI and the ratio ${\mathscr{R}}(G)$ in Eq. (16). These correlations are very close to the previous ones. Indeed, both $\kappa(G)$ and ${\mathscr{R}}(G)$ produce time series with very similar behaviors, making the two measures almost equivalent for application purposes. This fact can be explained by the presence of a large spectral gap in eigenvalues of the correlation matrices for each time window. In this case, it is possible to apply the approximate formula for the ratio in Eq. (20). This explains the almost-1 correlation between $\kappa(G)$ and ${\mathscr{R}}(G)$. We observe that the error between the exact value of the ratio in Eq. (19) and its approximate version in Eq. (20) is very small. Until October 2016, the mean relative error is $0.033\%$, with a maximum equal to $0.201\%$ in August, 2015. In the period from November 2016 to January 2020, the approximation error is on average $1.22\%$ with a max of $3.91\%$ in October 2017. The maximum is achieved in correspondence with a severe reduction of the spectral gap in 2017, see Fig. 5.2. We also observe that the network is perfectly balanced in the windows ending in February to May 2010, and starting at the end of 2008.
@@
Similarly, the correlations between the global balance and, for instance, the AMRI for $H=2$ are $\rho_{\rm Pearson}=0.5307341$ and $\rho_{\rm Spearman}=0.7685456$, on the period 2005-2016 and $\rho_{\rm Pearson}=0.4994806$ and $\rho_{\rm Spearman}=0.9441659$, on the period 2016-2020. The correlation between the global balance and the Cumulative Risk Fraction are $\rho_{\rm Pearson}=0.9096708$ and $\rho_{\rm Spearman}=0.9203074$, on the period 2005-2016 and $\rho_{\rm Pearson}=0.9297902$ and $\rho_{\rm Spearman}=0.991976$, on the period 2016-2020. To test the robustness of our findings, we repeated the exercise over many different subsets of 50 assets randomly chosen from the original index, obtaining similar results (see Fig. 5.3).
Figure 5.3: Time evolution of the global balance (black square dot line) and of the market rank indicator for $p=3$ and $H=2$ (blue kite dot line), $H=5$ (red triangle dot line) and $H=10$ (green circle dot line) for the period (a) 2005-2016 and (b) 2016-2020. The figures refer to a subset of $50$ randomly selected assets.
@@ 5.2 Effectiveness of the Global Balance as a Systemic Risk Measure
We perform now an empirical analysis to assess the effectiveness of the global balance in detecting systemic events. In order to identify systemic events, we refer to the definition in [54]. A systemic risk event occurs if the average return of all the $N$ securities falls below a threshold $\tau$ for $\Delta T$ consecutive days. We adopt a threshold $\tau=-0.01$ or $\tau=-0.005$ and sliding windows of $\Delta T=20$ days.
Figure 5.4: Panel (a): Average returns of the S&P500 database in the period 2005-2020. The average returns of each asset have been computed over a $20$-days-wide sliding window with $1$-day step. The blue line represents the mean over all the assets in the same window of the average returns. The green and the red lines represent the maximum and the minimum, respectively, of the average returns in the same window. Panel (b) focuses on the mean of the average returns and highlights crisis events when the line goes below the two possible threshold represented by the two horizontal dashed lines.
In Fig, 5.4, panel (a), we plot the average returns of the S&P500 database from January 4, 2005 to March 18, 2020. The average returns of each asset have been computed over a $20$-days-wide sliding window with $1$-day step. The blue line represents the mean of the average returns over all the $385$ assets in the same window. The green and the red lines represent the maximum and the minimum value, respectively, of the average returns in the same window. Panel (b) focuses on the mean of the average returns and highlights crisis events when the line goes below the two possible thresholds represented by the two horizontal dashed lines.
In this way, we can clearly identify systemic crisis events when the mean of the average returns over all the assets goes below $\tau$, for both $\tau=-0.005$ and $\tau=-0.010$. In particular, the crisis events below $\tau=-0.010$ ranges in the time intervals September-November 2008, February 2009, July 2011, October 2019.
We study the joint distribution of the average returns and of the global balance in the sliding windows. The range of the indicator $\kappa(G)$ has been divided into $n$ bins $B_{1},B_{2},\dots,B_{n}$ and the conditional distributions of the average returns $P\left(\langle X_{it}\rangle|\,\kappa(G)\in B_{k}\right)$ have been computed on each bin. Specifically, in a first experiment the interval $[0,1]$ is divided into $5$ bins of equal width while, in a second scenario, the five intervals are $(0,0.5]$, $(0.5,0.8]$, $(0.8,0.9]$, $(0.9,0.99]$ and $(0.99,1.00]$. The necessity of testing what happens with intervals of different lengths, very small when close to 1, is a peculiarity in the framework of systemic risk. Since systemic risk events are, hopefully, rare compared to the periods in which the market behaves normally, a systemic risk indicator should discriminate a few events within a multitude.
@@
Figure 5.5 displays the densities of the average returns conditioned to the values of $\kappa(G)$, obtained by using a Gaussian kernel smoothing for the whole dataset of $385$ assets (panels (a) and (b) for the weighted version of the correlation networks and panels (c) and (d) for the binary version). When $\kappa(G)$ increases, the densities move leftwards and flatten. More precisely, three effects can be recognized in relation to the increase of $\kappa(G)$: the average of the conditional distribution decreases; the standard deviation increases; the VaR [^5] The Value at Risk, $p$ VaR, measures the loss that will not be exceeded with a given confidence probability $p$ over a certain time horizon (see, for instance, [55]). We adopt $p=0.05$. of the average returns increases. Fig. 5.6 shows the boxplots of the same distributions conditioned on the two different partitions of the $[0,1]$ interval for the global balance in the weighted and binary version. The mean of the average returns within each interval decreases as $\kappa(G)$ increases, while the standard deviation increases. This is particularly noticeable by considering the extreme intervals in Fig. 5.6, panels (b) and (d), which refer to the events in the interval $\kappa(G)\in(0.99,1.00]$. The values of the mean of the conditional distributions in Fig. 5.6 are collected in Table 5.2. The means of the smoothed distributions in Fig. 5.5 are reported in the Supplementary Material, Sec. 5. Both are characterized by a strictly negative monotonic behavior as $\kappa(G)$ increases. Finally, we computed the VaR within each band. For instance, for the whole network in the weighted version and with intervals of equal width the VaR’s are: $0.02384863$ for $0<\kappa(G)\leq 0.2$; $0.0281525$, for $0.2<\kappa(G)\leq 0.4$; $0.03010301$, for $0.4<\kappa(G)\leq 0.6$; $0.03521984$, for $0.6<\kappa(G)\leq 0.8$; and $0.05100622$ for $0.8<\kappa(G)\leq 1.0$.
@@
This evidence shows that the global balance is a meaningful systemic risk measure. Indeed, a large value of $\kappa(G)$ signals that the market is experiencing worse performance and that the probability of big losses is high. We test the robustness of these numerical results on a database (ESX50) independent of the one analyzed here. We show the relative results in the Supplementary Material, Sec. 6.
6 Conclusion
This paper provides a bridge between the realm of balance in signed networks and that of systemic risk measures for asset correlation networks. The link between the two frameworks is built through the definition of a discrete diffusive process that drives the spread of the information on the network. On the one side, the process permits to derive the global balance indicator in an alternative way with respect to the standard combinatoric approach. On the other side, the steady state of the diffusive process is the solution of a linear system where the coefficients are the elements of the exponential matrix. Then the relation between the condition number of the exponential matrix and the global balance of the network follows in a straightforward way. The structural predictability of a network refers to the monotone behavior of the nodes in response to a perturbation at one or more nodes, and is equivalent to the numerical stability of a linear system. We prove interesting properties that shed light on the relationship between the global balance and the condition numbers of the signed and unsigned problems. We also introduce an almost complete characterization of signed networks that consider the number of possible odd/even positive/negative closed walks. By replacing the adjacency matrix with the correlation matrix of financial returns, we apply the global balance index to systemic risk detection. This step is natural considering that there exists a whole class of systemic risk measures proposed in the literature that are generalizations of the condition number of the correlation matrix. The application to two different databases of real financial data supports and confirms our idea. The global balance of the correlation network can be used as a systemic risk indicator. The main advantage of this approach is that it does not require a full rank correlation matrix. Consequently, the global balance of the correlation network can be calculated also when the number of observations is smaller than the number of the financial assets. This scenario is of great interest in the framework of systemic risk for two main reasons. First, ideally the number of securities constituting a good approximation of the whole financial index is supposed to be very large, and our methodology allows to deal with a large number of assets. Second, our approach allows to work with short estimation windows containing only more recent data, which are necessary to obtain a systemic risk measure with a good descriptive content and a potential predictive power. There is one further advantage that characterizes our proposal. The global balance indicator is naturally obtained as the aggregation of the local balance indices. Hence, these local indices measure the contribution of the single node in the correlation network to the global balance. While standard approaches to systemic risk usually measures the overall risk of the economic system, the approach based on the balance is basically local. Our future research will focus on the local balance to investigate the role of each asset to the general stability of the market.
Data availability Data will be made available on request. The code used to compute the indicators used in this article can be found in the following repository:
Acknowledgments We thank the editor and anonymous reviewers for helpful comments and suggestions. FDD acknowledges funding from Ministerio de Ciencia e Innovacion, Agencia Estatal de Investigacion Program for Units of Excellences Maria de Maeztu (CEX2021--001164-M///10.13039///501100011033), from MDM2017071120220170711202-2017-0711-20-2- 2017 - 0711 - 20 - 2 funded by MCIN///AEI///10.13039///
501100011033 and by FSE invierte en tu futuro, as well as project APASOS (No. PID2021 --122256NB--C22). RG and PU acknowledge financial support from the European Union – NextGenerationEU. Project PRIN 2022 “Networks:decomposition, clustering and community detection” code: 2022NAZ0365 - CUP H53D23002510006. PB and RG are members of GNAMPA-INdAM.
Supplementary material
1 Another possible choice of the coefficient $\alpha(k)$ in Eqs. (1) and (2)
We show here another analytically tractable choice for the coefficient $\alpha(k)$ in Eq. (2) of the main text. For simplicity, we only consider the limit $t\to\infty$. We assume throughout the derivation that the spectral radius of $\bf A$ is smaller than one. We begin with
${\bf G}={\bf I}+\sum_{j=1}^{\infty}\left(\prod_{k=0}^{j-1}\alpha(k)\right){\bf A}^{j},$
with
$\alpha(k)=\frac{(k+2)^{2}}{(k+2)^{2}-1}=\frac{(k+2)^{2}}{(k+1)(k+3)}.$
We first compute the product. Define
$P(j)=\prod_{k=0}^{j-1}\alpha(k)=\prod_{k=0}^{j-1}\frac{(k+2)^{2}}{(k+1)(k+3)}.$
Trivial algebraic manipulations show that
$P(j)=\frac{2\,((j+1)!)^{2}}{\,j!(j+2)!}=\frac{2(j+1)}{j+2}.$
Substituting $P(j)$ into $\bf G$ and doing some algebra, we get to:
${\bf G}=\textbf{I}+2\sum_{j=1}^{\infty}\left(1-\frac{1}{j+2}\right){\bf A}^{j}=\textbf{I}+2{\bf A}(\textbf{I}-{\bf A})^{-1}-2\sum_{j=1}^{\infty}\frac{{\bf A}^{j}}{j+2}.$
We evaluate the remaining sum:
$\sum_{j=1}^{\infty}\frac{{\bf A}^{j}}{j+2}={\bf A}^{-2}\sum_{j=3}^{\infty}\frac{{\bf A}^{j}}{j}={\bf A}^{-2}\Bigl{(}-\ln(\textbf{I}-{\bf A})-{\bf A}-\frac{{\bf A}^{2}}{2}\Bigr{)},$
where we have used the Taylor expansion of the logarithm. Substituting back into the expression for ${\bf G}$:
${\bf G}=2\textbf{I}+2{\bf A}(\textbf{I}-{\bf A})^{-1}+2{\bf A}^{-1}+2{\bf A}^{-2}\ln(\textbf{I}-{\bf A}).$
2 Proof of Proposition 1
Proposition 1.
Proof.
$=\frac{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}-W^{-}_{2k}\right)+\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}-W^{-}_{2k+1}\right)\right]}{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}+W^{-}_{2k}\right)+\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}+W^{-}_{2k+1}\right)\right]}\cdot$
$\quad\ \frac{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}-W^{-}_{2k}\right)-\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}-W^{-}_{2k+1}\right)\right]}{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}+W^{-}_{2k}\right)-\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}+W^{-}_{2k+1}\right)\right]}$
where $W^{\pm}_{\rm even}$ and $W^{\pm}_{\rm odd}$ denote the total penalized weight of positive and negative closed walks of even and odd length, respectively. For a balanced graph, by definition, there are no negative closed walks and $W^{-}_{\rm even}=W^{-}_{\rm odd}=0$. For an antibalanced graph, by definition, every even closed walk is positive and every odd closed walk is negative, so $W^{-}_{\rm even}=W^{+}_{\rm odd}=0$. The claim follows. ∎
3 Results for random correlation matrices
Proposition 2.
$\sum_{i=1}^{N}e^{\lambda_{i}}\sim I_{2}(2)e^{2}N\sim 5.090679N$
and
$\sum_{i=1}^{N}e^{-\lambda_{i}}\sim\frac{I_{0}(2)+I_{1}(2)}{e^{2}}N\sim 0.523778N$
Proof.
$M_{k}=\frac{1}{N}\sum_{i=1}\lambda_{i}^{k}\sim\binom{2k}{k}\frac{1}{k+1}=\frac{(2k)!}{k!(k+1)!}$
Then we have
$\begin{split}\sum_{i=1}^{N}e^{-\lambda_{i}}&=\sum_{i=1}^{N}\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{k!}\lambda_{i}^{k}=\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{k!}\sum_{i=1}^{N}\lambda_{i}^{k}\\ &=N\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{k!}M_{k}=N\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{k!}\frac{(2k)!}{k!(k+1)!}\\ &=N\sum_{k=0}^{\infty}(-1)^{k}\frac{(2k)!}{(k!)^{3}(k+1)}=N\frac{I_{0}(2)+I_{1}(2)}{e^{2}}\end{split}$
As a consequence, the condition number grows, under these assumptions, as $2.67N^{2}$, that is $\mathscr{K}(e^{-{\bf A}})\sim O(N^{2})$. A similar result can hardly be replicated for the absolute value matrix, i.e., for the network $|G|$. Nonetheless, we can investigate the statistical correlation between the ratio of the two condition numbers ${\mathscr{R}}({\bf A})=\mathscr{K}(e^{-{\bf A}})/\mathscr{K}(e^{-|{\bf A}|})$ and the global balance $\kappa(G)$. In particular, we generate random correlation matrices [^6] Random matrices are generated by randcorr package in R. of increasing size $N=5$, $N=10$, $N=20$, $N=50$, and we consider for each dimension 100 repetitions. In Fig. 3.1, we report the scatterplot of the values of ${\mathscr{R}}({\bf A})$ and $\kappa(G)$. As can be seen, the values of the two indicators correlate very strongly and increasingly as the network size, i.e. the number of assets involved, increases.
Figure 3.1: Increasing correlation between the ratio ${\mathscr{R}}({\bf A})$ and the global balance $\kappa(G)$ for (a) $N=5$, $\rho=0.8546633$, (b) $N=10$, $\rho=0.9693817$, (c) $N=20$, $\rho=0.9918888$, (d) $N=50$, $\rho=0.9984123$
4 Global balance and average correlation
In Fig. 4.1, we show the temporal evolution of the global balance $\kappa(G)$ and of the average correlation for the S&P500 dataset. In Fig. 4.1 we have split the entire interval into two sub-intervals, 2005-2016 in panel (a) and 2016-2020 in panel (b), in order to highlight with higher resolution the information carried by the indicators in the two different periods. In Fig. 4.2, we zoom in on the period 2008-2014, when the global balance undergoes small relative changes.
Figure 4.1: Global Balance (square dots black line) and Average Correlation (circular dots red lines) for the S&P500 daily returns with $\Delta T=400$ and $\Delta t=30$, for the period (a) 2005-2016 and (c) 2016-2020.
Figure 4.2: Global Balance (square dots black line) and Average Correlation (circular dots red lines) for the S&P500 daily returns with $\Delta T=400$ and $\Delta t=30$, zoom on the period 2008-2014.
In Figs. 4.3 and 4.4, we show the temporal evolution of the global balance $\kappa(G)$ and of the average correlation for the subset of $50$ assets in the S&P500 dataset, considered in the main text, again over the whole time interval and the partitioned time interval, respectively.
Figure 4.3: Global Balance (square dots black line) and Average Correlation (circular dots red lines), a sample of $50$ assets from S&P500 daily returns, $\Delta T=100$ and $\Delta t=30$, for the period 2005-2016.
Figure 4.4: Global Balance (square dots black line) and Average Correlation (circular dots red lines), a sample of $50$ assets from S&P500 daily returns, $\Delta T=100$ and $\Delta t=30$, for the period 2016-2020.
5 Mean of the smoothed conditional distributions of the average returns.
@@ 4 Global balance and average correlation
6 Eurostoxx 50 index analysis
We check the robustness of our results using a different database than the one used in the main text. We refer here to the constituents of the eurostoxx 50 index (ESX50). The cleaned dataset contains the daily returns of $42$ stocks from January 5, 2005 to September 17, 2020. We perform the test over windows of $45$ days. In Fig. 6.1 we plot the mean, maximum and minimum average returns of the constituents, panel (a), and a zoom on the mean, panel (b). Due to the larger size of the time windows, we detect less but more severe events. In this new European dataset, we specifically identify the systemic crisis associated with COVID-19 in January 2020 with a threshold $\tau=-0.010$. Fig. 6.2 and Table 6.1 illustrate the conditional smoothed distributions, the boxplots and the corresponding mean values for the weighted version of the correlation networks. We do not show the results for the binary network for the present example. The numerical results confirm the previous findings, reinforcing the interpretation of the global balance as an effective systemic risk indicator.